OpenStudy (anonymous):

Solve: |x + 3| = 3 A. {0, 6} B. {0, -6} C. {-6, 6} D. {-6}

OpenStudy (callisto):

|x + 3| = 3 x + 3 = 3 or x + 3 = -3 Can you solve it from here??

OpenStudy (anonymous):

yes thanks

OpenStudy (callisto):

welcome :)